Question 4 Quality Control As the number of electronic devices increases, so does the use of particular manufacturer produces batteries in lots of 100. In each lot, two of the batteries will be defective. The batteries are randomly packaged in groups of four batteries. What is the probability that all of the batteries in a package will not be defective? rechargeable batteries. A Question 5 Calculate the following a) C b) 17 P c) 1s C16 d) s P2 2. 1 Question 6 Athere is a 60% probabity of going to the dance, and a 40% of staying home And there is a 30% probability that will rain n Los Angeles Find the probabity of gong to the dance and clear
Answer #1

5)

a)6C3 =6!/(3!*(6-3)!)=6*5*4*3*2*1/(3*2*1*3*2*1)=20

b) 17P3 =17!/(17-3)! =17*16*15=4080

c)as factorial a negative number can not be calculated ; it is not applicable' another way to see this is that we can not choose 16 from 15 objects

d)3P2 =3!/(3-2)!=3*2*1=6

e)

8P1 =8!/(8-1)!=8

6)

probability=P(dance)*P(clear skies|dance)=0.6*(1-0.3)=0.42